Part A Constants What is the capacitance per unit length? Express your answer using two significant figures. A cylindrical capacitor has an inner conductor with a radius of 1.2 mm and an outer conductor with a radius of 3.2 mm. The two conductors are separated by vacuum, and the entire capacitor has a length of 3.2 m C/L Submit incorrect; Try Again; 2 attempts remaining Part B The potential of the inner conductor relative to that of the outer conductor is 360 mV. Find the charge (magnitude and sign) on both conductors. Enter your answers numerically separated by a comma. Express your answers using two significant figures. Enter , Gouter-
Answer #1

Radius of inner conductor of a cylindrical capacitor= a = 1.2 mm

Radius of outer conductor = b = 3.2 mm .

The two conductors are separated by vacuum.Length of entire capacitor =L= 3.2 m

The capacitance per unit length = C/L= 2πε/ln(b/a)

The capacitance per unit length = C/L= 55.60619*10-12/ln(3.2/1.2)

The capacitance per unit length = C/L= 55.60619*10-12/0.9808
The capacitance per unit length = C/L= 56.695 *10-12 F/m

b) The potential of the inner conductor relative to that of the outer conductor is 360 mV .

The charge (magnitude and sign) on both conductors= q =CV

q =( 56.695 *10-12 )*3.2*0.36

Charge on outer= q = + 6.531264*10-11 C

Charge on inner= q = - 6.531264*10-11 C